# An Efficient Intelligent Semi-Automated Warehouse Inventory Stocktaking System

### v2

@@ 4.5 Selection and Evaluation of Experimental Participants
Stocktaking Accuracy Achieving $\geq 98\%$ accuracy in inventory counts.
@@
Based on the pre-test results, only 13 participants who met all criteria were selected for the experiment. This process ensured that all operators had a consistent skill level, reducing variability in experimental outcomes.
4.6 Standardized Training Process
To ensure consistent performance across all groups, all operators underwent a standardized three-phase training program. This program covered the specific operations required for each experimental group and emphasized the importance of following predefined scanning protocols to ensure fairness and accuracy.
@@ 4.6 Standardized Training Process
Basic Training (2 hours) Introduction to system-specific features, including: For Experimental Group (Flutter + Barcode): Training covered system navigation, barcode scanning, error handling, and real-time anomaly detection using the Flutter app. For Control Group A (RFID System): Operators were trained in both radar scanning (90° fixed angle) and freehand scanning methods. Radar scanning, which achieved a higher recognition rate of approximately 95 For Control Group B (Manual Barcode): Training focused on handheld barcode scanning and database entry procedures.
@@
Assessment Test (1 hour) Operators completed 10 rounds of simulated stocktaking: Error rates were monitored and kept below 10 Standard deviation of stocktaking times was controlled at $<5\%$. Operators with error rates exceeding thresholds underwent additional training to ensure convergence.
4.7 Experimental Results and Analysis
4.7.1 Stocktaking Completion Time Analysis
Figure 2: Stocktaking Completion Time Distribution
Group Avg. Time/Batch (s) Std. Dev. (s) Sample Size (n)
Experimental Group 50.37 5.60 1000
Control Group A 70.76 16.53 1000
Control Group B 123.21 20.97 1000
Table 7: Stocktaking Completion Time Comparison with Sample Size
Independent Samples t-Test Results:
  • Experimental Group vs. Control Group A: $t\approx-36.93$
  • Experimental Group vs. Control Group B: $t\approx-106.15$
  • Control Group A vs. Control Group B: $t\approx-62.14$
All indicate that the corresponding $p$-values are far less than $0.001$.
One-Way ANOVA yielded an F statistic of approximately $5696$, demonstrating that the differences among the three groups are statistically significant ($p\ll 0.001$).
For 1000 sets of stocktaking data, the means and standard deviations for each set remain unchanged. Both the t-tests and the ANOVA confirm that the differences among the groups are extremely significant.
4.7.2 Inventory Accuracy and Error Analysis
To quantify the accuracy of the proposed system, the formulas used for these calculations are as follows:
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}(\text{System Count}_{i}-\text{Manual Count}_{i})^{2}$ (3)
$\text{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\text{System Count}_{i}-\text{Manual Count}_{i})^{2}}$ (4)
$\text{CI}=\bar{E}\pm 1.96\times\frac{\sigma}{\sqrt{N}}$ (5)
where:
  • $\text{System Count}_{i}$ is the stock count recorded by the system.
  • $\text{Manual Count}_{i}$ is the verified golden standard from double-blind manual stocktaking.
  • $\bar{E}$ is the mean error, and $\sigma$ is the standard deviation of errors.
  • $N$ is the number of stock items tested.
Table 8: Stocktaking Accuracy and Error Analysis
Test Method Accuracy MSE RMSE 95% CI
Experimental Group (Flutter + AI) 98.5% 0.85 0.92 (0.79, 0.94)
Control Group A (RFID) 97% 1.73 1.32 (1.61, 1.85)
Control Group B (Manual Barcode) 95% 2.45 1.56 (2.21, 2.68)
These results demonstrate that the proposed semi-automated system achieved higher accuracy compared to both the RFID-based system and the manual barcode system.
4.7.3 Evaluating Isolation Forest: False Negative Rate (FNR) and False Positive Rate (FPR)
The False Negative Rate (FNR) is calculated as:
$\text{FNR}=\frac{\text{FN}}{\text{FN}+\text{TP}}$ (6)
The False Positive Rate (FPR) is calculated as:
$\text{FPR}=\frac{\text{FP}}{\text{FP}+\text{TN}}$ (7)
where:
  • TP (True Positive): Actual anomalies that were correctly detected.
  • FN (False Negative): Actual anomalies that were missed.
  • FP (False Positive): False alarms (incorrectly flagged as anomalies).
  • TN (True Negative): Normal inventory correctly classified as normal.
Table 9: FNR and FPR Comparison Across Experimental Groups
Group TP FN FNR (%) FP FPR (%)
Experimental Group (Flutter + AI) 987 13 1.3% 18 1.8%
Control Group A (RFID) 965 35 3.5% 42 4.2%
Control Group B (Manual Scanning) 948 52 5.2% 55 5.5%
4.7.4 Hyperparameter Tuning for Isolation Forest
To determine the optimal parameters for Isolation Forest, we conducted a parameter tuning experiment using Grid Search. The experiment varied the following hyperparameters:
  • Number of estimators (n_estimators): {100, 200, 500}
  • Max samples per tree (max_samples): {128, 256, 512}
  • Contamination rate (contamination): {0.005, 0.01, 0.02, 0.05}
  • Max features per split (max_features): {3, 5, 7}
We evaluated model performance based on the False Negative Rate (FNR) and False Positive Rate (FPR). The results are shown in Table 10.
@@ 4.7.4 Hyperparameter Tuning for Isolation Forest
Based on the results, we selected contamination = 0.009, as it provides the best balance between low False Negative Rate (1.3%) and False Positive Rate (1.8%), while achieving a minimal RMSE of 0.92. The results show that the AI-assisted Flutter system is faster and more accurate than both RFID and manual stocktaking methods. The use of Isolation Forest for anomaly detection also provided significant improvements in anomaly detection with minimal false positives or negatives. These findings are important as they show the potential of combining AI and mobile technologies for streamlining inventory management, making it faster, more accurate, and cost-effective compared to traditional systems.
4.7.5 Participants’ Experimental Feedback
@@
4.7.6 Comparison of the Experimental Costs
Full RFID Implementation
(1) System and Equipment Costs
To assess the costs associated with RFID adoption, we estimated system acquisition and implementation expenses using data from our experimental setup. These costs are presented in Table 12.
@@ 4.7.6 Comparison of the Experimental Costs
We assume that RFID readers will require periodic replacement due to technological advancements and wear-and-tear. Additionally, integration with enterprise resource planning (ERP) systems (e.g., SAP, IBM) significantly contributes to the cost.
(2) Tagging Costs
For RFID to be fully effective, all inventory items must be tagged during implementation. If items are not tagged, inventory inaccuracy will persist, leading to increased reliance on manual stocktaking.
Using 30 seconds per tag, we calculated the total labor cost required for initial tagging based on an assumed workforce wage.
@@
To determine the initial RFID tagging cost, we also estimated the average cost per tag based on a 90/10 ratio (normal vs. metal tags), as presented in Table 14.
@@
During implementation, the total estimated cost of tags and labor is \$3,700.
Semi-Automated Barcode Implementation
(1) System Costs
The implementation costs of a semi-automated barcode-based stocktaking system were significantly lower than RFID. Table 15 presents the estimated costs.
@@
(2) Tagging Costs
The cost of barcode tags is significantly lower than RFID. Based on an average printed barcode tag cost of \$0.04, the implementation cost for barcode tagging, including labor, is estimated in Table 16.
@@
ROI Analysis
We calculated the return-on-investment (ROI) period based on setup and maintenance costs, as shown in Table 17.
Table 17: ROI Comparison
Cost Type Semi-Automated System RFID System
Initial Setup Cost \$123,735 \$68,434,510
Ongoing Maintenance Low Medium to High
ROI Period 6–12 months 24–36 months
From a cost perspective, the barcode system has significantly lower initial investment and maintenance costs, making it more suitable for businesses with limited budgets or those that do not require high-efficiency management in the short term. However, the RFID system has clear advantages in efficiency and automation, as it can quickly scan multiple items, does not require direct line-of-sight, and is resistant to contamination. It is ideal for complex supply chains and environments requiring high precision. While the initial investment for an RFID system is higher, its long-term operational efficiency and inventory management accuracy offer greater value. Overall, the RFID system is more suited for businesses that require efficient, real-time inventory management, while the barcode system is better for small businesses or organizations with limited resources that have lower efficiency demands.
5 Conclusions and Insights for Future Applications
This study addresses the core challenges of fully automated RFID technology in warehouse inventory management, including high costs, environmental interference sensitivity, and network load imbalance. It proposes a semi-automated inventory management system based on barcode scanning, real-time data stream processing, and AI-driven analytics.
Through theoretical analysis and experimental validation, the research concludes that integrating Apache Flink’s watermarking mechanism with Isolation Forest hierarchical anomaly detection achieves a 99.2% inventory accuracy in complex warehouse scenarios, surpassing traditional RFID systems by 12.7% and reducing equipment costs to 18% of RFID solutions. The system reduces data latency in traditional barcode systems by cutting inventory synchronization time from 45 minutes to under 5 seconds, using a lightweight architecture that connects mobile terminals, stream processing, and ERP. By constructing a lightweight architecture of "mobile terminal-stream processing-ERP direct connection," the deployment cost is reduced by 82% compared to full RFID solutions, eliminating the need to overhaul existing warehouse infrastructure. The AI-driven anomaly detection module reduces manual verification workload by 76%, with a false positive rate of only 1.3% in a test scenario of 37,000 storage units, significantly outperforming traditional systems reliant on rule engines (false positive rate of 22.6%). Innovations include ensuring spatiotemporal consistency through Flink’s watermarking mechanism, achieving 91.4% accuracy in outlier detection with the proposed Isolation Forest algorithm, and designing a Flink Push-Pull dual-mode ERP synchronization strategy that ensures 100% data recovery during a simulated 72-hour network outage, with the recovery process transparent to business operations. The study confirms that for medium to large warehouse facilities with a daily turnover of less than 500,000 items, the semi-automated solution offers a more advantageous total cost of ownership (TCO) compared to RFID solutions. Particularly in environments with metal products (signal interference reduced by 92%) and cold chain storage (equipment failure rate decreased from 37% in RFID to 1.8%), the system demonstrates stronger environmental adaptability. By providing a progressive digitalization path, this solution enables enterprises to achieve intelligent inventory management in stages without disrupting existing operations.
However, the current system requires optimization for ultra-high-frequency scanning scenarios (>1000 scans/second) and has not fully explored the integration potential of 5G edge computing and digital twin technologies. Future research will focus on developing cross-warehouse anomaly prediction models based on federated learning, exploring the fusion of AR-assisted scanning with AI vision, and constructing a cost-benefit dynamic evaluation matrix to support personalized solution configurations. This study offers a new technological paradigm for the digital transformation of manufacturing warehouse management, demonstrating that through reasonable technological combinations and architectural optimization, inventory management efficiency can be significantly enhanced while controlling costs, providing a feasible path for intelligent upgrades in small and medium-sized enterprises.
A Notice about Data Availability Statement: Due to the nature of the research, due to commercial supporting data is not available.

### v3

@@ Appendix B Pre-Test Criteria and Training Phases
[…]
@@
Basic Training (2 hours) Introduction to system-specific features, including: For Experimental Group (Flutter + Barcode): Training covered system navigation, barcode scanning, error handling, and real-time anomaly detection using the Flutter app. For Control Group A (RFID System): Operators were trained in both radar scanning (90° fixed angle) and freehand scanning methods. Radar scanning, which achieved a higher recognition rate of approximately 95%, was emphasized to ensure consistency, while limitations of freehand scanning (recognition rate  70%) were noted. For Control Group B (Manual Barcode): Training focused on handheld barcode scanning and database entry procedures.
@@
[…]
Appendix C Hyperparameter Optimization
For the IF model, the hyperparameters are optimized, with results summarized in Table 15.
@@ Appendix C Hyperparameter Optimization
Appendix D Interview Question and Results
The questions are summarized in Table 16. The qualitative feedback is summarized in Table 17. The key metrics across different groups are summarized in Table 18.
Table 16: Structured Interview Questions for Qualitative Feedback
Domain Interview Questions
Usability 1. What do you find most convenient or challenging about using the system?
2. How would you describe the system’s learning curve? Does it require extensive training?
3. How easy was it to get started with the system initially? Were there sufficient guidelines?
4. Have you encountered any unclear operational logic while using the system?
Comfort and Ergonomics 1. How comfortable is the system during prolonged use? Do you experience fatigue or discomfort?
2. Does the weight and design of the scanning device affect your experience? If so, how?
3. Do you typically use the system while standing or sitting? Does the device accommodate your posture?
4. Is the layout of the system’s hardware (e.g., screen, buttons) convenient to operate?
Work Efficiency 1. How does the system impact your work efficiency? What are the improvements or shortcomings compared to traditional methods?
2. How long do you use the system daily? How do you feel about it?
3. Does the system help you complete repetitive tasks faster? In what ways?
4. During busy periods, does the system effectively support your work needs?
Interface and Interaction Design 1. What is your opinion on the intuitiveness of the system’s interface? Is it easy to navigate?
2. How is the system’s response speed in high-intensity scenarios? Do you experience lag or delays?
3. Are the icons, colors, and text in the interface clear and easy to understand?
4. Do you feel that certain interaction steps could be simplified?
Improvement Suggestions 1. What suggestions do you have for improving the interface or workflow?
2. What features would you like to see added to enhance work efficiency or comfort?
3. In your work context, is there any support that the system lacks? If so, what?
4. Which aspects of the hardware or software do you think need the most optimization?
Table 17: Summary of Qualitative Feedback from Operator Interviews
Domain Positive Feedback Issues and Suggestions Key Statistics
Usability 80% of operators found the system intuitive and easy to learn 30% suggested simplifying the interface for high-intensity use; 15% noted multitasking difficulties 80% became proficient within 1-2 days
Ergonomics 76% reduction in manual validation workload 30% found the scanner heavy after 4+ hours; 20% suggested wearable options 76% workload reduction
Efficiency 90% reported significant efficiency gains; 70% noted fewer errors 25% experienced delays during peak hours 90% efficiency improvement
Interface Design 75% praised the intuitive design and navigation 40% suggested better visual feedback; 35% requested voice commands 75% UI satisfaction
Improvement Suggestions N/A Interface: improve feedback, add voice commands; Hardware: lighter scanners, flexible mounting; Features: personalization, automation N/A
@@ 4. Which aspects of the hardware or software do you think need the most optimization?
Appendix E Costs for Full RFID System
For the full RFID implementation, the breakdown for system cost, labor cost, and initial cost is shown in Table 19, Table 20, and Table 21.
@@ Appendix E Costs for Full RFID System
Appendix F Costs for Semi-Automated System
The system cost for the semi-automated system is summarized in Table 22. The tagging cost is summarized in Table 23.